# Don’t be Fooled: The Misinformation Effect of Explanations in Human-AI Collaboration

### v1

@@ Abstract.
Across various applications, humans increasingly use black-box artificial intelligence (AI) systems without insight into these systems’ reasoning. To counter this opacity, explainable AI (XAI) methods e enhanced transparency and interpretability. While recent studies have explored how XAI affects human-AI collaboration, few have examined the potential pitfalls caused by incorrect explanations. The implications for humans can be far-reaching but have not been explored extensively. To investigate this, we ran a study (n=160) on AI-assisted decision-making in which humans were supported by XAI. Our findings reveal a misinformation effect when incorrect explanations accompany correct AI advice with implications post-collaboration. This effect causes humans to infer flawed reasoning strategies, hindering task execution and demonstrating impaired procedural knowledge. Additionally, incorrect explanations compromise human-AI team-performance during collaboration. With our work, we contribute to HCI by providing empirical evidence for the negative consequences of incorrect explanations on humans post-collaboration and outlining guidelines for designers of AI.
@@ 1. Introduction
Imagine you are studying for an art history exam and must know how to distinguish two architectural styles. You seek advice from an online artificial intelligence (AI) assistant for explanations to differentiate the two styles. Despite being plausible, the AI’s explanation is incorrect. Yet you learn from these incorrect explanations, and as a consequence, your understanding of architectural styles is impaired! You fail your exam.
@@
Recent rapid technological advancements, in particular generative AI with the popular example of ChatGPT, have significantly bolstered both the capabilities and the adoption of AI (Bommasani et al., 2021; Dwivedi et al., 2023). However, as AI technologies get more sophisticated, the complexity and opacity of supported decision-making processes increase as well, presenting new challenges to ensure that these systems are still transparent and interpretable for humans (Rudin et al., 2022). This increasing complexity necessitates a deeper understanding of the underlying factors impacting human-AI interaction, especially in scenarios where human oversight is critical (Amershi et al., 2019; Sterz et al., 2024). Regulatory frameworks, such as the EU AI Act, mandate human oversight to ensure AI systems operate ethically, legally, and safely (Commission, 2021). This underscores the significance of HCI research to develop methods and tools that facilitate effective collaboration between humans and AI (Shneiderman, 2020). Ensuring that AI is not only accurate but also explainable is vital for fostering appropriate reliance and complementary team performance (Schemmer et al., 2023c; Hemmer et al., 2023).
The need for explainability in AI has been widely acknowledged, leading to substantial advancements in both research and practice (Silva et al., 2023; Arrieta et al., 2020). In the financial sector, for instance, eXplainable AI (XAI) is used in credit scoring and fraud detection, targeting improved auditing, regulatory compliance, and user trust (Cirqueira et al., 2021).
@@
Despite these advancements, the XAI literature has barely picked up on an important potential pitfall in human-AI collaboration scenarios: incorrect explanations (Morrison et al., 2024; Lakkaraju and Bastani, 2020) for accurate AI advice may impair humans post-collaboration as they compromise their procedural knowledge—their ability to perform specific tasks on their own— and their reasoning—their ability to derive conclusions based on their understanding. Exploring the effect of incorrect explanations on procedural knowledge and reasoning is crucial, as regulations (Commission, 2021) or performance expectations (Hemmer et al., 2023) might require humans to complement AI knowledge with their own knowledge and, therefore, maintain the ability to perform tasks on a superior performance level. This is particularly critical in high-stakes domains like healthcare, finance, and legal settings, where flawed decisions can have severe consequences (Rudin, 2019). Additionally, in scenarios where AI advice is first provided and then removed, like in AI-based teaching settings (Spitzer et al., 2023, 2024b), maintaining humans’ procedural knowledge is vital.
@@
Understanding the impact of incorrect explanations in AI-assisted decision-making even when the AI advice is correct is essential for designing more effective collaborations between humans and AI. By identifying how and why incorrect explanations impact humans not only during, but in particular post-collaboration with AI, we can develop effective strategies to mitigate these effects (Chen et al., 2022; Lakkaraju and Bastani, 2020; Ehsan et al., 2019; Ribeiro et al., 2016). Thus, we ask the following research questions (RQs):
@@
To address these questions, we first synthesize how to measure the impact of incorrect explanations in AI-assisted decision-making based on previous research. We pre-registered our hypotheses on AsPredicted.org111We attached an anonymized copy of the pre-registration to the submission of this article.. Through an online study with 160 participants, we examine how incorrect explanations influence humans’ procedural knowledge (RQ 1) and reasoning (RQ 2) in a task to classify architectural styles of buildings. We provide participants with different types of AI support: no support at all, AI advice without explanations, AI advice with correct explanations, and AI advice with incorrect explanations. We measure the effects on their task performance during and after collaboration to derive the impact on their procedural knowledge and qualitatively analyze their understanding of the task to derive their reasoning ability. Additionally, we analyze the effect on the human-AI team performance (RQ 3).
The findings of our study reveal a misinformation effect in AI-assisted decision-making: incorrect explanations significantly impair humans, resulting in a notable decline in their procedural knowledge once they have to perform the task autonomously. We also find that humans’ reasoning is impaired when they receive incorrect explanations. In fact, a negative effect during the collaboration can also be substantiated in our study: human-AI teams perform worse when the AI provides incorrect explanations, curtailing the complementary benefits of this collaboration. These findings underscore the potential dangers of incorrect explanations and highlight the importance of developing robust and reliable explanatory support for humans.
In summary, our study makes several contributions to the field: first, it fills a critical gap in the literature by examining the impact of incorrect explanations on humans in AI-assisted decision-making. Second, it identifies and measures the misinformation effect within AI explanations and outlines its negative repercussions on humans—a decline in procedural knowledge and reasoning. Lastly, our research extends existing knowledge on the interplay between AI and human decision-makers, providing insights into the hazards of explanations on the human-AI team. Overall, this paper sheds light on the potential pitfalls of incorrect explanations and their implications for humans post-collaboration. With this exploratory work, we hope to contribute to the development of concepts and hypotheses helping to advance theoretical knowledge in XAI and AI-assisted decision-making as well as to further advance the design of more effective AI-based decision support systems.
@@ 2. Background
With the rapid advancement of AI and its integration into diverse decision-making processes, XAI has emerged as a critical technique for enhancing transparency and assistance to help decision-makers understand AI’s reasoning (Alufaisan et al., 2021). Especially with applications based on generative AI (like Open AI’s ChatGPT or Anthropic’s Claude), explanations are being generated in natural language that provide human-understandable support for the AI’s response (Singh et al., 2024; Zytek et al., 2024). Previous research on human-AI collaboration has predominantly focused on how AI explanations influence human decision-making (Subramanian et al., 2024; Schemmer et al., 2022). For example, Hemmer et al. (2021) examine the factors that impact human-AI team performance, suggesting that XAI can foster complementary collaboration between humans and AI. Albeit the positive effects, studies also highlight how XAI can impair the collaboration between humans and AI (Miller, 2019; Ehsan et al., 2021; Binns, 2018). However, the understanding of negative consequences of XAI is still limited (Mohseni et al., 2021), especially empirical evidence for the negative impact of incorrect explanations is missing. In Table 1, we sort recent works in AI-assisted decision-making according to the correctness of AI advice and explanations. The table shows the under-explored topic of incorrect explanations in HCI. We review recent literature that highlights the risks and limitations of collaborative settings between humans and AI, motivating the need for further exploration of how incorrect explanations impair humans.
The table shows in a two times two matrix how previous work contributed to correct/incorrect AI advice and correct/incorrect explanations. Correct AI explanations Incorrect AI explanations Correct AI advice (Schemmer et al., 2022; Hemmer et al., 2023; Schoeffer et al., 2024; Lai and Tan, 2019; Adhikari et al., 2019; Hase and Bansal, 2020; Ribeiro et al., 2018; van der Waa et al., 2021; Zhang et al., 2020; Yeung et al., 2020) (Morrison et al., 2024; Cabitza et al., 2024) Incorrect AI advice (Sadeghi et al., 2024; Schemmer et al., 2023c; Kim et al., 2023; Cau et al., 2023; Buçinca et al., 2021; Alufaisan et al., 2021; Ehrlich et al., 2011; Chen et al., 2023; Cabitza et al., 2023; Kocielnik et al., 2019) (Papenmeier et al., 2019; Morrison et al., 2024; Lakkaraju and Bastani, 2020)
@@ 2.1. Incorrect AI Advice in Human-AI Collaboration
Research in HCI has explored the effects of incorrect AI advice on human-AI collaboration (Kocielnik et al., 2019; Bansal et al., 2019). Kocielnik et al. (2019) propose three strategies for managing user expectations in scenarios involving incorrect AI advice: accuracy indicators, example-based explanations, and performance control. Through a study involving an AI-based scheduling assistant prone to errors, they demonstrate these techniques’ effectiveness in maintaining user satisfaction and acceptance despite occasional incorrect recommendations. Their findings also reveal that the specific nature of AI errors can significantly influence user perception and trust.
Several recent studies have investigated how programmers collaborate with Copilot, an AI programming assistant that is not always accurate (e.g., (Vasconcelos et al., 2023; Barke et al., 2023; Dakhel et al., 2023)). For instance, Vasconcelos et al. (2023) focus on an AI supporting code completion. As the AI can make errors in producing the code, they convey the uncertainty of Copilot’s outputs by highlighting code sections that are most likely to be edited by the programmer. Their findings reveal that this approach helps programmers to arrive at solutions more quickly. Additionally, Dakhel et al. (2023) conclude that, while GitHub Copilot is a valuable tool for expert programmers, non-expert programmers should exercise caution when using it due to the potential for errors. Further research in the HCI domain has delved into the broader implications of interacting with incorrect AI. Barke et al. (2023) examine how different levels of expertise among programmers influence their interaction with AI-generated code. Their results indicate that while expert programmers can effectively navigate and correct AI-generated errors, novice programmers often struggle, leading to decreased efficiency and increased frustration. Complementary studies have also explored how incorrect AI recommendations affect user trust and decision-making in various contexts. Yin et al. (2019) investigate user strategies for coping with incorrect AI advice in everyday tasks, discovering that users often develop heuristics to cross-verify AI recommendations with other information sources. Moreover, Yang et al. (2019) analyze the role of trust in human-AI collaboration, particularly in scenarios where AI occasionally provides incorrect advice. Their findings suggest that initial trust can facilitate smoother interactions, but that repeated exposure to incorrect advice significantly erodes trust.
@@ 2.2. Incorrect Explanations in Human-AI Collaboration
Next to the AI advice, the understanding in HCI of how incorrect explanations can have an impact in AI-assisted decision-making is limited. Only few studies investigate the effect of explanations’ incorrectness (Morrison et al., 2024; Cabitza et al., 2024; Lakkaraju and Bastani, 2020; Papenmeier et al., 2019). A recent study in HCI shows that not only incorrect advice but also incorrect explanations have the potential to deceive decision-makers (Morrison et al., 2024). The authors of Morrison et al. (2024) explore the negative impacts of such incorrect explanations on humans’ decision-making behavior. They extend the conceptualization of Schemmer et al. (2023c) by the explanation dimension and explore, in a bird classification study, how the correctness of explanations impacts humans’ reliance on AI. They show that incorrect explanations can deceive decision-makers, leading to an inappropriate reliance behavior. Cabitza et al. (2024) explore the effects of explanations in a logic puzzle task. They also show that incoherent explanations can mislead humans, resulting in an inappropriate reliance behavior. In another study, Papenmeier et al. (2019) run a study in AI-assisted decision-making with incorrect explanations. They study how the explanations affect humans’ trust in determining the publication of Tweets. Similarly, Lakkaraju and Bastani (2020) find that incorrect explanations can affect humans’ trust in AI by investigating their effects in law and criminal justice use-cases.
These studies collectively underscore the complex dynamics of human-AI interaction, especially how incorrect advice and incorrect explanations affect humans’ reliance behavior on AI. However, the HCI field lacks a deeper understanding of the effects of such impaired collaboration scenarios on humans themselves. Especially for scenarios in which not the advice but the explanation for the decision maker—intended to foster interpretability—is incorrect. Studies like Morrison et al. (2024) build a promising starting point to inform HCI researchers and practitioners of the downsides of incorrect explanations. However, we still do not know anything about the impact of these AI shortcomings on humans’ ability to perform the tasks autonomously (procedural knowledge) and to conclude about the underlying domain (reasoning) post collaboration.

### v2

@@ Abstract.
Across various applications, humans increasingly use black-box artificial intelligence (AI) systems without insight into these systems’ reasoning. To counter this opacity, explainable AI (XAI) methods promise enhanced transparency and interpretability. While recent studies have explored how XAI affects human-AI collaboration, few have examined the potential pitfalls caused by incorrect explanations. The implications for humans can be far-reaching but have not been explored extensively. To investigate this, we ran a study (n=160) on AI-assisted decision-making in which humans were supported by XAI. Our findings reveal a misinformation effect when incorrect explanations accompany correct AI advice with implications post-collaboration. This effect causes humans to infer flawed reasoning strategies, hindering task execution and demonstrating impaired procedural knowledge. Additionally, incorrect explanations compromise human-AI team-performance during collaboration. With our work, we contribute to HCI by providing empirical evidence for the negative consequences of incorrect explanations on humans post-collaboration and outlining guidelines for designers of AI.
@@ 1. Introduction
Imagine you are studying for an art history exam and must know how to distinguish two architectural styles. You seek advice from an online artificial intelligence (AI) assistant for explanations to differentiate the two styles. Despite being plausible and providing you with correct architectural styles, the AI’s explanation is incorrect. Yet, you learn from these incorrect explanations, and as a consequence, your understanding and capability to distinguish the architectural styles is impaired! You fail your exam.
@@
Recent technological advancements have significantly bolstered both the capabilities and the adoption of AI (Bommasani et al., 2021; Dwivedi et al., 2023). For instance, large language models (LLMs) can assist in high-stakes classification tasks, such as categorizing legal cases based on case descriptions, by providing initial advice along with an explanation, even when users have no prior expertise in legal analysis (Ma et al., 2024). In the medical field, AI has been used to accurately classify retinal disorders, leading to early diagnosis and providing critical support to ophthalmologists (Mukhtorov et al., 2023). However, as these AI technologies become more sophisticated, especially with the use of generative AI, the complexity and opacity of supported decision-making processes increase. This presents new challenges in ensuring that these systems are still transparent and interpretable for humans (Rudin et al., 2022). This increasing complexity necessitates a deeper understanding of the underlying factors that impact human-AI interaction, especially in scenarios where human oversight is critical (Amershi et al., 2019; Sterz et al., 2024; Cabitza et al., 2023a). Regulatory frameworks, such as the EU AI Act (Commission, 2021), mandate human oversight to ensure AI systems operate ethically, legally, and safely. This underscores the importance of human-computer interaction (HCI) research to develop methods and tools that facilitate effective collaboration between humans and AI (Shneiderman, 2020). Ensuring that AI is not only accurate but also explainable is vital to foster appropriate reliance and complementary team performance (Schemmer et al., 2023b; Hemmer et al., 2023). The need for explainability in AI has been widely acknowledged, leading to substantial advancements in both research and practice (Silva et al., 2023; Arrieta et al., 2020). In the financial sector, for instance, eXplainable AI (XAI) is used in credit scoring and fraud detection, targeting improved auditing, regulatory compliance, and user trust (Cirqueira et al., 2021).
@@
Despite these advancements, the XAI literature has barely picked up on an important potential pitfall in human-AI collaboration scenarios: incorrect explanations (Morrison et al., 2024; Lakkaraju and Bastani, 2020; Kayser et al., 2024) for accurate AI advice. These incorrect explanations may impair humans post-collaboration as they compromise their procedural knowledge—their ability to perform specific tasks on their own— and their reasoning—their ability to derive conclusions based on their understanding. Exploring the effect of incorrect explanations on procedural knowledge and reasoning is crucial, as regulations (Commission, 2021) or performance expectations (Hemmer et al., 2023) might require humans to complement AI capabilities with their domain knowledge and, therefore, maintain the ability to perform tasks on a superior performance level. This is particularly critical in high-stakes domains like healthcare, finance, and legal settings, where flawed decisions can have severe consequences (Rudin, 2019). The preservation of human procedural knowledge becomes even more essential in scenarios where AI advice is initially provided but later removed, such as in AI-based teaching settings (Spitzer et al., 2023, 2024b), as humans need to maintain their ability to perform tasks independently when AI support is unavailable.
@@
Understanding the impact of incorrect explanations in AI-assisted decision-making, even when the AI advice is correct, is essential for designing more effective collaborations between humans and AI. By identifying how and why incorrect explanations impact humans not only during, but in particular post-collaboration with AI, we can develop effective strategies to mitigate these effects (Chen et al., 2022; Lakkaraju and Bastani, 2020; Ehsan et al., 2019; Ribeiro et al., 2016). Thus, we ask the following research questions (RQs):
@@
To address these questions, we first synthesize how to measure the impact of incorrect explanations in AI-assisted decision-making based on previous research. We pre-registered our hypotheses on AsPredicted.org111We attached an anonymized copy of the pre-registration to the submission of this article.. Through an online study with 160 participants, we examine how incorrect explanations influence humans’ procedural knowledge (RQ 1) and reasoning (RQ 2) in a task to classify architectural styles of buildings. We chose this classification task because AI, and especially LLMs, are increasingly being used nowadays not only to generate content but also to support human reasoning and decision-making (Hadi et al., 2024). Thus, they provide advice to humans in scenarios where no prior knowledge is needed (e.g., patients seeking medical advice) to help them interpret and draw conclusions from complex data. Such settings remain crucial as they provide a controlled environment (clear ground truth and measurable outcomes) to study how humans are impacted by incorrect explanations. In this task, we provide participants with different types of AI support: no support at all, AI advice without explanations, AI advice with correct explanations, and AI advice with incorrect explanations. We measure the effects on their task performance during and after collaboration to derive the impact on their procedural knowledge and qualitatively analyze their understanding of the task to derive their reasoning ability. Additionally, we analyze the effect on the human-AI team performance (RQ 3).
The findings of our study reveal a misinformation effect in AI-assisted decision-making: incorrect explanations significantly impair humans, resulting in a notable decline in their procedural knowledge once they have to perform the task autonomously. We also find that humans’ reasoning is impaired when they receive incorrect explanations. In fact, we also observe a negative effect during the collaboration in our study: human-AI teams perform worse when the AI provides incorrect explanations, curtailing the complementary benefits of this collaboration. These findings underscore the potential dangers of incorrect explanations and highlight the importance of developing robust and reliable explanatory support for humans.
In summary, our study makes several contributions to the field: first, it fills a critical gap in the literature by examining the impact of incorrect explanations on humans in AI-assisted decision-making. Second, it identifies and measures the misinformation effect within AI explanations and outlines its negative repercussions on humans—a decline in procedural knowledge and reasoning. Lastly, our research extends existing knowledge on the interplay between AI and human decision-makers, providing insights into the hazards of explanations on the human-AI team. Overall, this paper sheds light on the potential pitfalls of incorrect explanations and their implications for humans post-collaboration. With this exploratory work, we hope to contribute to the development of concepts and hypotheses helping to advance theoretical knowledge in XAI and AI-assisted decision-making as well as to advance further the design of more effective AI-based decision support systems.
@@ 2. Background
With the rapid advancement of AI and its integration into diverse decision-making processes, XAI has emerged as a critical technique for enhancing transparency and assistance to help decision-makers understand AI’s reasoning (Alufaisan et al., 2021; Cabitza et al., 2024b). Especially with applications based on generative AI (like Open AI’s ChatGPT or Anthropic’s Claude), explanations are being generated in natural language that provide human-understandable support for the AI’s response (Singh et al., 2024; Zytek et al., 2024). Previous research on human-AI collaboration has predominantly focused on how AI explanations influence human decision-making (Subramanian et al., 2024; Schemmer et al., 2022). For example, Hemmer et al. (2021) examine the factors that impact human-AI team performance, suggesting that XAI can foster complementary collaboration between humans and AI. Albeit the positive effects, studies also highlight how XAI can impair the collaboration between humans and AI (Miller, 2019; Ehsan et al., 2021; Binns, 2018). However, the understanding of the negative consequences of XAI is still limited (Mohseni et al., 2021), especially empirical evidence for the negative impact of incorrect explanations is missing. In Table 1, we sort recent works in AI-assisted decision-making according to the correctness of AI advice and explanations. The table shows the under-explored topic of incorrect explanations in HCI. We review recent literature highlighting the risks and limitations of collaborative settings between humans and AI, motivating the need to explore further how incorrect explanations impair humans.
The table shows in a two times two matrix how previous work contributed to correct/incorrect AI advice and correct/incorrect explanations. Correct AI explanations Incorrect AI explanations Correct AI advice (Schemmer et al., 2022; Hemmer et al., 2023; Schoeffer et al., 2024; Lai and Tan, 2019; Adhikari et al., 2019; Hase and Bansal, 2020; Ribeiro et al., 2018; van der Waa et al., 2021; Zhang et al., 2020; Yeung et al., 2020) (Morrison et al., 2024; Cabitza et al., 2024a; Kayser et al., 2024) Incorrect AI advice (Sadeghi et al., 2024; Vicente and Matute, 2023; Kim et al., 2023; Cau et al., 2023; Buçinca et al., 2021; Alufaisan et al., 2021; Ehrlich et al., 2011; Chen et al., 2023; Cabitza et al., 2023b; Kocielnik et al., 2019) (Papenmeier et al., 2019; Morrison et al., 2024; Lakkaraju and Bastani, 2020; Kayser et al., 2024)
@@ 2.1. Incorrect AI Advice in Human-AI Collaboration
Research in HCI has explored the effects of incorrect AI advice on human-AI collaboration (Kocielnik et al., 2019; Bansal et al., 2019; Yin et al., 2019; Vicente and Matute, 2023). Kocielnik et al. (2019) investigate how such incorrect AI advice influences user satisfaction and acceptance. In their study, they demonstrate how interventions like accuracy indicators or performance control can maintain users’ perception and trust even when the AI-based scheduling assistant provides wrong advice.
Next to scheduling assistants, several recent studies have investigated how programmers collaborate with Copilot, an AI programming assistant that is not always accurate (e.g., (Vasconcelos et al., 2023a; Barke et al., 2023; Dakhel et al., 2023)). For instance, Vasconcelos et al. (2023a) focus on an AI supporting code completion. As the AI can make errors in producing the code, they convey the uncertainty of Copilot’s outputs by highlighting code sections that are most likely to be edited by the programmer. Their findings reveal that this approach helps programmers to arrive at solutions more quickly. Additionally, Dakhel et al. (2023) conclude that, while GitHub Copilot is a valuable tool for expert programmers, non-expert programmers should exercise caution when using it due to the potential for errors. Other studies further examine how different levels of expertise among programmers influence their interaction with AI-generated code (Barke et al., 2023). Their findings suggest that while expert programmers can effectively navigate and correct AI-generated errors, novice programmers often struggle, leading to decreased efficiency and increased frustration. While previous research shows the impact of incorrect AI advice on humans with different levels of prior domain knowledge, Vicente and Matute (2023) run a study on medical diagnosis to investigate how humans are influenced during collaboration with an AI that provides erroneous advice. They show that humans inherit the same errors as the AI, thus impairing their ability to conduct the task themselves. Building on this, previous research also shows how non-expert users are influenced by AI systems providing inaccurate advice (Schemmer et al., 2023a). They find that users with limited domain expertise were prone to overreliance on the AI’s outputs, resulting in decision errors. Contrarily, (Vasconcelos et al., 2023b) show that explanations can reduce overreliance by introducing a cost-benefit framework. They outline that explanations to be effective need to diminish the costs of verifying AI advice.
This review of prior research illustrates the potential negative impact of incorrect AI advice on decision outcomes and highlights how people’s existing knowledge shapes their ability to interpret and respond to AI output. Building on these findings, the next section explores the role of incorrect explanations and how they influence humans’ decision-making when supported by AI.
@@ 2.2. Incorrect Explanations in Human-AI Collaboration
Next to the AI advice, the understanding of how incorrect explanations can impact AI-assisted decision-making is limited. Only a few studies investigate the effect of explanations’ incorrectness (Morrison et al., 2024; Cabitza et al., 2024a; Lakkaraju and Bastani, 2020; Papenmeier et al., 2019). A recent study in HCI shows that not only incorrect advice but also incorrect explanations have the potential to deceive decision-makers (Morrison et al., 2024). Morrison et al. (2024) explore the negative impacts of incorrect explanations on humans’ decision-making behavior. They extend the conceptualization of Schemmer et al. (2023b) by the explanation dimension and explore, in a bird classification study, how the correctness of explanations impacts humans’ reliance on AI. They show that incorrect explanations can deceive decision-makers who possess no prior domain knowledge, leading to inappropriate reliance behavior. Cabitza et al. (2024a) explore the effects of explanations in a logic puzzle task. They also show that if advice and the accompanying explanation do not align, humans are misled ultimately resulting in inappropriate reliance behavior. Papenmeier et al. (2019) conduct another study in AI-assisted decision-making with incorrect explanations. They study how explanations affect humans’ trust in identifying offensive tweets. Similarly, Lakkaraju and Bastani (2020) find that incorrect explanations can affect humans’ trust in AI by investigating their effects in law and criminal justice use cases.
These studies collectively underscore the complex dynamics of human-AI interaction, especially how incorrect advice and incorrect explanations affect humans’ reliance behavior on AI. However, the HCI field lacks a deeper understanding of the effects of such impaired collaboration scenarios on humans themselves. Especially for scenarios in which not the advice but the explanation for the decision-maker—intended to foster interpretability—is incorrect. Studies like Morrison et al. (2024) and Cabitza et al. (2024a) build a promising starting point to inform HCI researchers and practitioners of the downsides of incorrect explanations. However, we still do not know anything about the impact of these AI shortcomings on humans’ ability to perform the tasks autonomously (procedural knowledge) and to conclude about the underlying domain (reasoning) post-collaboration.